OpenStudy (ivancsc1996):

URGENT!!!!!! I need someone to help me go through the steps of solving for v in this equation:

OpenStudy (ivancsc1996):

\[L=\frac{ v _{c}^{2}\cos \theta sen \theta -v _{c}\cos \theta \sqrt{v _{c}^{2}\sin ^{2}\theta-2g(h _{C}-h _{D})} }{ g }\]

OpenStudy (anonymous):

what is sen(theta)?

OpenStudy (ivancsc1996):

Like the numerical value of sin theta? It is just the factor that transforms v into a component in the y direction.

OpenStudy (anonymous):

is it supposed to be sin(theta)

OpenStudy (ivancsc1996):

Ahhh yeah. I speak spanish so for me sine is seno, it slipped sorry. Yes it is sin(theta)

OpenStudy (anonymous):

no problem.

OpenStudy (ivancsc1996):

@satellite73

OpenStudy (anonymous):

first multiply both sides by g

OpenStudy (anonymous):

then subtract the term vc^2cos(theta)*(sin(theta)) from both sides.

OpenStudy (anonymous):

then divide both sides by -vc*cos(theta)

OpenStudy (anonymous):

then square both sides.

OpenStudy (anonymous):

give me one sec. ill post a picture.

OpenStudy (ivancsc1996):

I did what you said but I got stuck again at:\[\frac{ L ^{2}g ^{2} }{ 2\cos(\theta) }=v _{c}^{2}(\cos(\theta)g(h _{C}-h _{D})+Lgsin (\theta)-v _{c}^{2}\cos(\theta)sen ^{2}(\theta))\]

OpenStudy (anonymous):

at the end it should be + or - square root.

OpenStudy (ivancsc1996):

Thank you!!!!!!!!! That was great.

OpenStudy (anonymous):

you're welcome

OpenStudy (ivancsc1996):

But where did the cos squared go in the eight step?